Pamela borrows an amount of money for emergency house repairs. The interest on the loan is compounded


quarterly. After four years the debt accumulated by R7 980,00 to the amount of R32 923,00. The yearly


interest rate, expressed as a percentage and rounded to two decimal places, at which the money was borrowed,


is


[1] 1,75%.


[2] 8,00%.


[3] 9,26%.


[4] 7,00%

1
Expert's answer
2022-04-19T00:23:43-0400

According to compound interest formula

"R=P*(1+{\frac i n})^{t*n}\implies i =n*(({\frac R P})^{{\frac 1 {tn}}}-1)" , where P - initial amount, i - yearly interest rate, n - number of payments per year, t - number of years

So, in the given case we have

"i=4(({\frac {32923} {7980}})^{{\frac 1 {4*4}}}-1)=0.37" = 37% yearly